OpenStudy (anonymous):

Can somebody show me how to change an equation that's in standard form to slope and y-intercept form???

OpenStudy (anonymous):

Post the problem

OpenStudy (anonymous):

Right now Im trying to change 4x-y=-1.

OpenStudy (anonymous):

Alright, to get slope intercept form you just wanna isolate y on one side of the equation. So first you wanna subtract 4x from both sides, right?

OpenStudy (anonymous):

Ok so then itd be y=-1 - 4x?

OpenStudy (anonymous):

What happened to the negative in front of the y? Also, since you want the form y=mx+b just subtract the x after the equal sign -y=-4x-1 now you should do what to finish solving for y?

OpenStudy (anonymous):

The negative on teh side with y cancels out the negative on teh other side to make the yintercept form of the equation y=-4x +1?

OpenStudy (anonymous):

close, You dont just divide the 1 by -1, you divide everything so it would change both signs y=4x+1

OpenStudy (anonymous):

Ok.Can you help me change this one to please???-x-y=x-5

OpenStudy (anonymous):

Alright, what do you think you should do first?

OpenStudy (anonymous):

Add -x to each side.

OpenStudy (anonymous):

Yes add x to both sides, so what does that give you?

OpenStudy (anonymous):

y=2x-5

OpenStudy (anonymous):

right except the y is still negative -y=2x-5 now one more step

OpenStudy (anonymous):

y=2x+5

OpenStudy (anonymous):

Again, both are being divided by -1 so it changes both signs y=-2x+5

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

If you had -x+y=x-5...would the yintercept form of that equation be y=-2x-5 or did I do that wrong?

OpenStudy (anonymous):

Close, all you need to do in this one is add an x to both sides x+x=2x so y=2x-5

OpenStudy (anonymous):

Ok thanks again! :)

OpenStudy (anonymous):

If you have the equation 2x-y=-5...would the yintercept form be y=2x-5?? @CalebBeavers

OpenStudy (anonymous):

Nope, do it in steps 1: subtract 2x from both sides 2:divide both sides by -1 -y=-2x-5 y=2x+5

OpenStudy (anonymous):

Ok Can you help with this one...-2x-y=-1.I know I need to add -2x to both sides first.